Operations Report – March 8th

Crew 223 Operations Report 08/03/2020

SOL: 7

Name of person filing report: Luc FORTIN

Non-nominal systems: Nothing to report
Notes on non-nominal systems: Nothing to report

Saturday generator:
Off at 6:00 am
On at 8:57 pm SOC was 86%

Sunday generator:
Off at 7:30 am
On at 9:00 pm SOC was XX%

Diesel Reading – 68%
Station Propane Reading – 75%
Ethanol Free Gasoline – 5.5 gallons

Water (loft tank): 47 gallons
Water Meter: 149748.04 units
Water (static tank) – 280 gallons
Static to Loft Pump used – yes
Water in GreenHab – 135.2 gallons
Water in ScienceDome: 0 gallons

Toilet tank emptied: yes

Perseverance rover used: yes
Hours: 150.2
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Spirit rover: no
Hours: 134.1
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes.
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Curiosity rover used: yes
Hours: 139.5
Beginning charge: 100%
Ending charge: 96%
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (Honda, 1, 2 and 3)

Reason for use: I started the 4 ATVs during the engineer EVA for battery check. I had trouble starting the Honda ATV (same as sol 4). It sounds like a flat battery. We will try again to start it during the EVA tomorrow, and if not starting we will use the jumper cables or bring back the battery to charge it in the RAM.

Oil Added? no

ATV Fuel Used: Nothing to report

Hours the ATVs were Used today: Nothing to report.

Notes on ATVs: Nothing to report.

HabCar used and why, where? The HabCar is parked at the MDRS Campus.

CrewCar used and why, where? The CrewCar is in Grand Junction.

General notes and comments: Nothing to report

Summary of internet: All ok.
Summary of suits and radios: Suits 7, 10 were used during the EVA.
Summary of Hab operations: Nothing to report.
Summary of GreenHab operations: Nothing to report.
Summary of ScienceDome operations: Nothing to report.
Summary of RAM operations: Nothing to report.
Summary of any observatory issues: Nothing to report
Summary of health and safety issues: Nothing to report

Questions, concerns, and requests to Mission Support: Nothing to report